# Design principles of the cytotoxic CD8+ T-cell response

### v2

@@ Abstract
Cytotoxic T lymphocytes eliminate infected or malignant cells, safeguarding surrounding tissues. Although experimental and systems-immunology studies have cataloged many molecular and cellular actors involved in an immune response, the design principles governing how the speed and magnitude of T-cell responses emerge from cellular decision-making remain elusive. Here, we recast the T-cell response as a feedback-controlled program, wherein the rates of activation, proliferation, differentiation and death are regulated through antigenic, pro- and anti-inflammatory cues. By exploring a broad class of feedback-controller designs as potential immune programs, we demonstrate how the speed and magnitude of T-cell responses emerge from optimizing signal-feedback to protect against diverse infection settings. We recover an inherent trade-off: infection clearance at the cost of immunopathology. We show how this trade-off is encoded into the logic of T-cell responses by hierarchical sensitivity to different immune signals. Notably, we find that designs that balance harm from acute infections and autoimmunity produce immune responses consistent with experimentally observed patterns of T-cell effector expansion in mice. Extending our model to immune-based T-cell therapies for cancer tumors, we identify a trade-off between the affinity for tumor antigens (“quality”) and the abundance (“quantity”) of infused T-cells necessary for effective treatment. Finally, we show how therapeutic efficacy can be improved by targeted genetic perturbations to T-cells. Our findings offer a unified control-logic for cytotoxic T-cell responses and point to specific regulatory programs that can be engineered for more robust T-cell therapies.
@@ I Introduction
We then ask how alternative immunological objectives give rise to distinct effector programs. First, we parameterize a broad class of feedback-controller designs, with varied signal sensitivities and integration rules. Next, by systematically sampling designs, we optimize for rapid pathogen clearance while minimizing immunopathology. We find that optimal designs lie on a Pareto front: further reductions in infection-induced harm entail increased immunopathology. Accounting for the imperative to separate infection—microbial or malignant—from transient inflammatory events (e.g., tissue damage) [25, 26, 27, 28, 29], we show how a trade-off between pathogen-clearance and immunopathology (toxicity) may be encoded into the logic of T-cell responses. Optimizing these competing objectives yields effector programs consistent with experiments with vertebrate organisms. Finally, applying our framework to a minimal tumor-growth model, we demonstrate how modifying the endogenously encoded effector program can improve the effectiveness of T-cell cancer therapies, pointing to signal-processing adjustments—and thus candidate genetic engineering—that could enhance T-cell-based cancer immunotherapy.
@@
Figure 1: CD8+ T-cell response to infections. (A) The schematic shows how upon infection, susceptible cells $S$ become infected II at rate $b_{I}$. Infected cells provide an antigen signal $\sigma_{\text{Ag}}\propto I$, and die at rate $d_{I}$, which triggers an infection harm signal $\sigma_{\text{inf}}$. In response to infection, naïve CD8+ T-cells (with initial clone size of $N_{\text{lin.}}$) are activated $N\to N^{*}$, and can then differentiate into effector or memory fates (Fig. S1). Effector cells can proliferate and kill infected and bystander cells, generating a response-harm signal $\sigma_{\text{res}}$, and finally become terminal or differentiate into memory. Consequently, T-cell transitions are modulated by the threes signals ($\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}}$). Infection dynamics $(S,I)$ are modeled by ODEs coupled to a stochastic dynamics of the T-cell response. Memory transitions (gray arrows) are neglected in our model. (B) Representative trajectories of susceptible, infected, effector, and memory CD8+ T-cells are shown over time. (C) Each T-cell transition has a rate $r_{i}=r^{\text{max}}g_{i}(\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}};\vec{\psi},\ell)$, modulated by a sigmoidal function with baseline $\ell$ and signal sensitivities $\vec{\psi}$ (eqs. 6, 7). The regulatory parameters define the design of the effector program $\vec{\theta}$ (eq. 8). (D) We assess design performance over diverse infections sampled from a broad range of reproduction numbers $R_{0}$ (eq. 2) and immunogenicities $\kappa$ (eq. 3). For each infection, we can find an optimal design $\vec{\theta}^{(\mathrm{min})}$ that minimizes total cell loss (harm). Harm decomposes into infection harm $H_{\text{inf.}}$ (cells killed by the pathogen plus infected cells eliminated by effectors), and toxicity $H_{\text{toxic.}}$ (healthy bystanders killed by effectors), highlighting an intrinsic trade-off (eq. 9). Simulation parameters in (B): $S_{\text{max}}=10^{7},N_{\text{lin.}}=100,K_{I}=10^{-3}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},b_{I}=2.0\times 10^{-7}\text{day}^{-1}\penalty 10000\ \text{per cell},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}},\vec{\theta}=(\psi_{Ag}=2,\psi_{I}=2,\psi_{E}=-0.25,\ell_{N\to N^{*}}=-0.75,\ell_{N^{*}\to E}=-0.75,\ell_{E\to\emptyset}=0.75$).
Model
To characterize the immune response to an infection, we model the large-population dynamics of susceptible and infected cells in a host as deterministic processes, coupled to a stochastic, time-inhomogeneous birth-death process for the CD8+ T-cell response (Fig. 1A, and SI Appendix for details).
@@
We model a population of healthy, susceptible cells $S$, initially consisting of $S_{\text{max}}$ identical cells. Pathogen exposure seeds an infected subpopulation $I(t)$, which grows via encounters with susceptible cells at rate $b_{I}$, and declines by infected-cell death at rate $d_{I}$, triggering an immune response (Figs. 1A, S1). The dynamics of the susceptible and infected cells follow,
where $d_{E}$ is the maximal clearance rate of infected/susceptible cells by effector T-cells $E$, and $K_{I}$ and $K_{S}$ are recognition thresholds, quantifying how readily effector cells can recognize infected and susceptible cells, respectively, and consequently kill them; a larger $K_{I}$ or $K_{S}$ implies more effector cells are needed for recognition [44, 45, 46]. The functional form of the last term in these equations can be motivated by considering competitive binding of T-cells to infected and healthy cells (SI Appendix).
@@
measuring the expected number of secondary infected cells generated by a single infected cell in an otherwise fully susceptible cell population. This definition mirrors the standard epidemiological reproduction number [47, 48]. We model the T-cell repertoire as initially consisting of a number of identical naïve T-cells, and assume $K_{S}=S_{\text{max}}\geq K_{I}$ so that infected cells are typically easier to recognize than susceptible cells. These recognition thresholds determine the immunogenicity of an infection, which we define as,
@@
where $h_{\text{inf}}(t)$ represents the number of cells killed directly by the pathogen (infection), and $h_{E,I}(t)$ and $h_{E,S}(t)$ denote the number of infected and healthy cells, respectively, killed by the effector T-cell response, amounting to the total damage by the response $h_{\text{res}}(t)$.
@@
We simplify this picture by partitioning non-antigen cues into (i) infection-induced and (ii) response-induced harms, whose accumulated effects trigger infection and response harm-induced signals, $\sigma_{\text{inf}}(t)$ and $\sigma_{\text{res}}(t)$, respectively. Antigens presented by cells also serve as a signal; we take the infected-cell burden $I(t)$ as a proxy for the instantaneous antigen signal $\sigma_{\text{Ag}}(t)$. Together, expressed by their recognition thresholds, $K_{I}$ for antigens and $K_{H}$ for non-antigens (harm), the three signals available to T-cells follow,
The time-integrated harm-induced signals are exponentially discounted at rate $d_{H}$, reflecting the decay of cytokines in the absence of ongoing stimulation (SI Appendix).
By disentangling these sources of (potentially redundant) signals, we probe how CD8+ T-cells use these cues to make fate decisions. We do not pre-assign any of the signals to a pro- or anti-inflammatory role but allow optimization goals to determine their roles across diverse infection scenarios.
@@
These response processes map onto partially distinct genetic programs: the MYC pathway drives activation and proliferation [17]; the Wnt-Tcf7 pathway regulates effector versus memory fate decisions [51, 16]; and the Bcl-2-family of proteins control contraction of the effector cells via apoptosis [17].
@@
where the maximal rate $r^{\text{max}}$ is capped by the cell-cycle time [13, 21, 49], and $g_{i}$ is a monotonic response function of signal input, chosen from the modified Monod-Wyman-Changeux functions (Hill-type family) [52, 53]:
@@
Figure 2: Design parameters shape T-cell response and infection clearance. (A) Varying the design parameters influences effector trajectories. (B) Trajectories are shown for the population size of effectors (top) and infected cells (bottom), measured in units of maximum number of susceptible cells $S_{\text{max}}$, for designs that vary a single baseline transition rate $g_{0,i}$ (color) indicated atop each panel. All other design parameters are fixed at: fast activation, $g_{0,N\to N^{*}}=0.95$; fast differentiation, $g_{0,N^{*}\to E}=0.95$; slow contraction, $g_{0,E\to\emptyset}=0.05$; and zero signal feedback, $\psi_{\text{Ag}}=\psi_{\text{inf}}=\psi_{\text{res}}=0$. Trajectories reflecting the impact of varying signal sensitivities $\psi$’s are shown in SI Appendix, Fig. S2. (C) Peak effector count, measured in units of maximum number of susceptible cells $S_{\text{max}}$, is shown by varying each of the three baseline transition rates (colors), while keeping the rest of the parameters fixed (similar to (B)). Parameters are the same as Fig. 1, except for $K_{I}=10^{-2}S_{\text{max}}$.
@@
Tracking all clinical sequelae of an infection is rarely feasible, but tissue integrity degrades as cells are lost. Cumulative cell death correlates with macroscopic readouts such as weight change, pathogen burden, and temperature dynamics [59, 60]. We therefore adopt a cell-economics objective in which harm is measured by the total number of cells lost over the course of infection.
We evaluate the performance of a given design $\vec{\theta}$ by simulating the response dynamics (Fig. 1D) and computing the total cell death (or inflicted harm) as
@@
where $H_{\text{inf.}}(\vec{\theta})$ aggregates infection-related harm (cells killed by the pathogen, and infected cells eliminated by effectors), $H_{\text{toxic.}}(\vec{\theta})$ captures immunopathology (healthy bystander cells killed by effectors), over the time window $T$ (SI Appendix, Table S1). This single, interpretable metric balances clearance benefits against collateral damage, enabling principled comparison across designs and infections (Fig. 1D).
@@
Figure 3: Toxicity-infection harm trade-off across T-cell regulatory designs. (A) Density plots of toxicity $H_{\text{toxic.}}$ versus infection harm $H_{\text{inf.}}$ (eq. 9)—measured in units of $S_{\text{max}}$—for $4.8\times 10^{6}$ T-cell regulatory designs, $\vec{\theta}$, uniformly grid-sampled across four infection scenarios (panels; columns: $\kappa=5.6$ (left), $\kappa=178$ (right); rows: $R_{0}=2.5$ (bottom), $R_{0}=5.0$ (top)). Designs are sampled uniformly across signal sensitivities $\psi\,\text{'s}\in[-3,3]$ and baseline rates $\ell\,\text{'s}\in[-3,3]$ (eq. 8). Infection scenarios are sampled by changing $K_{I}$ and $b_{I}$, while holding $S_{\max}$, $d_{I}$, and $K_{S}$ fixed. Black points denote no-feedback designs ($\vec{\psi}=0$), and the large magenta circle in each panel marks the design that minimizes total harm, $\vec{\theta}^{(\mathrm{min})}$. (B) For many different infection scenarios $(\kappa,R_{0})$, we identify an ensemble of near-optimal designs $\Theta(\kappa,R_{0})$, whose total harm is within $10^{-2}\times S_{\text{max}}$ of the minimum sampled total harm (as in the magenta point in A). By computing the ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ for each infection scenario, and using these averages to cluster infections, we identify four infection “archetypes” (colors): I–IV, as indicated (SI Appendix). (C) Toxicity vs. infection harm associated with optimal designs for each archetype is shown. Large dots represent cluster averages and small transparent dots represent individual infections, sampled in (B). (D) Bars indicate the ensemble-mean of design parameters, $\langle\vec{\theta}\rangle_{\Theta^{(i)}}$, computed by identifying the ensemble of designs $\Theta^{(i)}$} which minimize cluster-averaged total harm for archetype $i$ to within $10^{-2}\times S_{\text{max}}$ of the minimum sampled cluster-averaged total harm (similar to (B)). Error bars denote the standard deviation within this ensemble. All design variables are normalized by their respective bounds ($\psi^{\max}=3$, $\ell^{\max}=3$). Unless otherwise noted, simulation parameters: $S_{\text{max}}=10^{7},N_{\text{lin.}}=10^{-5}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}.$
Results

### v3

@@ Abstract
Cytotoxic T lymphocytes eliminate infected or malignant cells, safeguarding surrounding tissues. Although experimental and systems-immunology studies have cataloged many molecular and cellular actors involved in an immune response, the design principles governing how the speed and magnitude of T-cell responses emerge from cellular decision-making remain elusive. Here, we recast T-cell response as a feedback-controlled program, wherein the rates of activation, proliferation, differentiation and death are regulated through antigenic, pro- and anti-inflammatory cues. By exploring a broad class of feedback-controller designs as potential immune programs, we demonstrate how the speed and magnitude of T-cell responses emerge from optimizing signal-feedback to protect against diverse infection settings. We recover an inherent trade-off: infection clearance at the cost of immunopathology. We show how this trade-off is encoded into the logic of T-cell responses by hierarchical sensitivity to different immune signals. Notably, we find the designs that balance the harm from acute infections and autoimmunity produce immune responses consistent with the experimentally observed patterns of T-cell effector expansion in mice. Extending our model to immune-based T-cell therapies for cancer tumors, we quantify the trade-off between the affinity for tumor antigens (“quality”) and the abundance (“quantity”) of infused T-cells necessary for effective treatment. Finally, we show how therapeutic efficacy can be improved by targeted genetic perturbations to T-cells. Our findings offer a unified control-logic for cytotoxic T-cell responses and point to specific regulatory programs that can be engineered for more robust T-cell therapies.
@@ I Introduction
We then ask how alternative immunological contexts require distinct effector programs. First, we parameterize a broad class of feedback-controller designs, with varied signal sensitivities and integration rules. Next, by systematically sampling designs, we optimize for rapid pathogen clearance while minimizing immunopathology. We find that optimal designs lie on a Pareto front: further reductions in infection-induced harm entail increased immunopathology. Accounting for the imperative to separate infection—microbial or malignant—from transient inflammatory events (e.g., tissue damage) [25, 26, 27, 28, 29], we show how a trade-off between pathogen-clearance and immunopathology (toxicity) may be encoded into the logic of T-cell responses. Optimizing these competing objectives yields effector programs consistent with experiments with vertebrate organisms. Finally, applying our framework to a minimal tumor-growth model, we demonstrate how modifying the endogenously encoded effector program can improve the effectiveness of T-cell cancer therapies, pointing to signal-processing adjustments through genetic engineering that could enhance T-cell-based cancer immunotherapy.
@@
Figure 1: CD8+ T-cell response to infections. (A) The schematic shows how upon infection, susceptible cells $S$ become infected II at rate $b_{I}$. Infected cells provide an antigen signal $\sigma_{\text{Ag}}\propto I$, and die at rate $d_{I}$, which triggers an infection harm signal $\sigma_{\text{inf}}$. In response to infection, naïve CD8+ T-cells (with initial clone size of $N_{\text{lin.}}$) are activated $N\to N^{*}$, and can then differentiate into effector or memory fates (Fig. S1). Effector cells can proliferate and kill infected and bystander cells, generating a response-harm signal $\sigma_{\text{res}}$, and finally become terminal or differentiate into memory. Consequently, T-cell transitions are modulated by the threes signals ($\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}}$). Infection dynamics $(S,I)$ are modeled by ODEs coupled to stochastic dynamics of the T-cell response. Memory transitions (gray arrows) are neglected in our model. (B) Representative trajectories of susceptible, infected, effector, and memory CD8+ T-cells are shown over time. Cell counts are reported in units of the maximum number of susceptible cells $S_{\text{max}}$. (C) Each T-cell transition has a rate $r_{i}=r^{\text{max}}g_{i}(\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}};\vec{\psi},\ell)$, modulated by a sigmoidal function with baseline $\ell$ and signal sensitivities $\vec{\psi}$ (eqs. 4, 5). The regulatory parameters define the design of the effector program $\vec{\theta}$ (eq. 6). (D) We assess design performance over diverse infections sampled from a broad range of reproduction numbers $R_{0}$ (eq. 1) and immunogenicities $\kappa$ (eq. 2). For each infection, we can find an optimal design $\vec{\theta}^{(\mathrm{min})}$ that minimizes total cell loss (harm). Harm decomposes into infection harm $H_{\text{inf.}}$ (cells killed by the pathogen plus infected cells eliminated by effectors), and toxicity $H_{\text{toxic.}}$ (healthy bystanders killed by effectors), highlighting an intrinsic trade-off (eq. 7). Simulation parameters in (B): $S_{\text{max}}=10^{7},N_{\text{lin.}}=100,K_{I}=10^{-3}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},b_{I}=2.0\times 10^{-7}\text{day}^{-1}\penalty 10000\ \text{per cell},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}},T=30\penalty 10000\ \text{days},\vec{\theta}=(\psi_{Ag}=2,\psi_{I}=2,\psi_{E}=-0.25,\ell_{N\to N^{*}}=-0.75,\ell_{N^{*}\to E}=-0.75,\ell_{E\to\emptyset}=0.75$).
II Model
To characterize the immune response to an infection, we model the large-population dynamics of susceptible and infected cells in a host as deterministic processes, coupled to a stochastic, time-inhomogeneous birth-death process for the CD8+ T-cell response (Fig. 1A, B and SI Appendix).
@@ II Model
We model a population of healthy, susceptible cells $S(t)$. Pathogen exposure seeds an infected subpopulation $I(t)$, which grows over time $t$ via encounters with an otherwise homeostatically stable susceptible cell population, and declines through infected-cell death, triggering an immune response (Figs. 1A, B S1). Activated effector T-cells $E(t)$ can kill both infected and susceptible cells, with clearance governed by recognition thresholds $K_{I}$ and $K_{S}$, respectively, which quantify how readily effector cells recognize each cellular subpopulation; higher thresholds indicate that more effector cells are needed for recognition and killing. We model this dynamics as a deterministic birth-death process with competitive interaction of T-cells with infected and healthy cells; see Materials and Methods and the SI Appendix.
@@
where $b_{I}/d_{I}$ is the ratio of birth to death rates for infected cells, and $S_{\text{max}}$ is the maximum number of susceptible cells in a tissue (Materials and Methods). $R_{0}$ measures the expected number of secondary infected cells generated by a single infected cell in an otherwise fully susceptible cell population. This definition mirrors the standard epidemiological reproduction number [44, 45].
We model the T-cell repertoire as initially consisting of a number of identical naïve T-cells, and assume that the T-cell recognition threshold for infected cells is lower than that for susceptible cells ($K_{I}\leq K_{S}$), such that T-cells more readily respond to infected cells. Accordingly, we define the immunogenicity of an infection as,
@@
where $h_{\text{inf}}(t)$ represents the number of cells killed directly by the pathogen (infection), and $h_{E,I}(t)$ and $h_{E,S}(t)$ denote the number of infected and healthy cells, respectively, killed by the effector T-cell response, amounting to the total damage from the response $h_{\text{res}}(t)$.
@@
We simplify this picture by partitioning non-antigen cues into (i) infection-induced and (ii) response-induced harms, whose accumulated effects trigger infection and response harm-induced signals, $\sigma_{\text{inf}}(t)$ and $\sigma_{\text{res}}(t)$, respectively. Presented antigens on APCs serve as an additional cue. We take the infected-cell burden $I(t)$ as a proxy for the instantaneous antigen signal $\sigma_{\text{Ag}}(t)$, which together with the other two signals, modulate the cellular decision-making processes in Fig. 1A; see Materials and Methods and SI Appendix for details. By disentangling these sources of (potentially redundant) signals, we probe how CD8+ T-cells use these cues to make fate decisions. We do not pre-assign any signals to a pro- or anti-inflammatory role but allow optimization goals to determine their influence across diverse infection scenarios.
We should emphasize that CD8+ responses do not occur in isolation: the innate immune system, regulatory T-cells (Tregs), helper T-cells, and B-cells can all directly or indirectly shape CD8+ dynamics. While these interactions are not modeled explicitly, we interpret the harm and antigen signals as coarse-grained proxies for inputs from other immune compartments, capturing their net pro- and anti-inflammatory effects. In the same vein, our immunogenicity parameter $\kappa$ (eq. 2) may also subsume regulatory constraints, e.g. through Treg-mediated suppression that limits effector responses to self-like antigens, which in our framework correspond to low-immunogenicity targets. Incorporating these additional components explicitly would be an important direction for future work, enabling the study of richer feedbacks and more granular immune regulation and dynamics.
@@
These response processes map onto partially distinct genetic programs: the MYC pathway drives activation and proliferation [17]; the Wnt-Tcf7 pathway regulates effector versus memory fate decisions [48, 16]; contraction can be attributed to effector cell apoptosis controlled by the Bcl-2-family of proteins, or the differentiation of effectors into effector-memory cells [16]. The latter is regulated by the balance of T-bet/Blimp-1 (effector) versus Bcl-6/Eomes (memory) programs, as well as metabolic pathways such as mTORC1 [49]. To reduce model degeneracy, we do not distinguish among these distinct biological sources of contraction in our analysis.
@@
where the maximal rate $r^{\text{max}}$ is capped by the cell-cycle speed [13, 21, 46], and $g_{i}$ is a monotonic response function of signal input, chosen from the modified Monod-Wyman-Changeux functions (Hill-type family) [50, 51]:
@@
Figure 2: Design parameters shape T-cell response and infection clearance. (A) Varying the design parameters influences effector trajectories. (B) Trajectories are shown for the population size of effectors (top) and infected cells (bottom) for designs that vary a single baseline transition rate $g_{0,i}$ (color) indicated atop each panel. Both quantities are measured in units of the maximum number of susceptible cells, $S_{\text{max}}$, which is a convenient reference scale for different cell sub-populations. All other design parameters are fixed at: fast activation, $g_{0,N\to N^{*}}=0.95$; fast differentiation, $g_{0,N^{*}\to E}=0.95$; slow contraction, $g_{0,E\to\emptyset}=0.05$; and zero signal feedback, $\psi_{\text{Ag}}=\psi_{\text{inf}}=\psi_{\text{res}}=0$. Trajectories reflecting the impact of varying signal sensitivities $\psi$’s are shown in SI Appendix, Fig. S2. (C) Peak effector count, measured in units of maximum number of susceptible cells $S_{\text{max}}$, is shown by varying each of the three baseline transition rates (colors), while keeping the rest of the parameters fixed (similar to (B)). Solid lines show the mean of effector counts across $n=5$ simulation replicates and shaded envelopes indicate one standard deviation (Materials and Methods). Parameters are the same as Fig. 1, except for $K_{I}=10^{-2}S_{\text{max}}$.
@@
Tracking all clinical sequelae of an infection is rarely feasible, but tissue integrity degrades as cells are lost. Cumulative cell death correlates with macroscopic readouts such as weight change, pathogen burden, and temperature dynamics [57, 58]. We therefore adopt a cell-economics objective in which harm is measured by the total number of cells lost over the course of infection. We evaluate the performance of a given design $\vec{\theta}$ by simulating the response dynamics (Fig. 1D) and computing the total cell death (or inflicted harm) as
@@
where $H_{\text{inf.}}(\vec{\theta})$ aggregates infection-related harm (cells killed by the pathogen, and infected cells eliminated by effectors), $H_{\text{toxic.}}(\vec{\theta})$ captures immunopathology (healthy bystander cells killed by effectors), over the fixed time window $T$. In our main simulations, we set $T=30$ days to reflect the typical timescale of acute infections and the associated immune response (SI Appendix, Table S1), but we also assess the sensitivity of our results to variation in $T$ (Materials and Methods). This single, interpretable metric of harm (eq. 7) balances clearance benefits against collateral damage, enabling principled comparison across designs and infections (Fig. 1D).
@@
Figure 3: Toxicity-infection harm trade-off across T-cell regulatory designs. (A) Density plots of toxicity $H_{\text{toxic.}}$ versus infection harm $H_{\text{inf.}}$ (eq. 7)—measured in units of the maximum number of susceptible cells $S_{\text{max}}$—for $4.8\times 10^{6}$ T-cell regulatory designs, $\vec{\theta}$, uniformly grid-sampled across four infection scenarios (panels; columns: $\kappa=5.6$ (left), $\kappa=178$ (right); rows: $R_{0}=2.5$ (bottom), $R_{0}=5.0$ (top)). Designs are sampled uniformly across signal sensitivities $\psi\,\text{'s}\in[-3,3]$ and baseline rates $\ell\,\text{'s}\in[-3,3]$ (eq. 6). Infection scenarios are sampled by changing $K_{I}$ and $b_{I}$, while holding $S_{\max}$, $d_{I}$, and $K_{S}$ fixed. Black points denote no-feedback designs ($\vec{\psi}=0$), and the large magenta circle in each panel marks the design that minimizes total harm, $\vec{\theta}^{(\mathrm{min})}$. (B) For many different infection scenarios $(\kappa,R_{0})$, we identify an ensemble of near-optimal designs $\Theta(\kappa,R_{0})$, whose total harm is within $10^{-2}\times S_{\text{max}}$ of the minimum sampled total harm (as in the magenta point in A). By computing the ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ for each infection scenario, and using these averages to cluster infections, we identify three pathogen “archetypes” (colors): I–III, as indicated (SI Appendix). (C) Toxicity vs. infection harm associated with the optimal harm-minimizing designs for each pathogen archetype in (B) is shown. Large points indicate archetype means; small semi-transparent points show individual immune challenges sampled in (B). Reducing infection harm (horizontal axis) comes at the expense of increased toxicity (vertical axis). (D) Bars indicate the ensemble-mean of design parameters, $\langle\vec{\theta}\rangle_{\Theta^{(i)}}$, computed by identifying the ensemble of designs $\Theta^{(i)}$} which minimize cluster-averaged total harm for each archetype within $10^{-2}\times S_{\text{max}}$ of the minimum sampled cluster-averaged total harm (similar to (B)). Error bars denote the standard deviation within this ensemble. All design variables are normalized by their respective bounds ($\psi^{\max}=3$, $\ell^{\max}=3$). Unless otherwise noted, results represent the average of $n=5$ simulation replicates with parameters: $S_{\text{max}}=10^{7},N_{\text{lin.}}=10^{-5}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}.$
III Results